OpenStudy (anonymous):

6x^2-13xy+6b^2 Factor completely and explain steps

OpenStudy (anonymous):

6x^2-13xy+6b^2 =6x^2-9xy-4xy+6y^2 =3x(2x-3y)-2y(2x-3y) =(3x-2y)(2x-3y)

OpenStudy (phi):

is the last term really 6 b^2 ? or is it 6 y^2 ?

OpenStudy (anonymous):

it's 6y^2. Sorry for the typo.

OpenStudy (anonymous):

First get the product of a and c and then find the factors of 36 will give you the sum of -13 product of a & c 6 * 6 = 36 Now find factors of 36 that gives us the sum of -13 -9 + -4 = -13 Now we solve by grouping. 6x^2-13xy+6b^2 Take out the -13xy and replace it with our two numbers -9xy and -4xy 6x^2-9xy-4xy+6y^2 Now Group (6x^2-9xy) (-4xy+6y^2) Find the factor for the first group (6x^2-9xy)/3x 3x(2x-3y) Find the factor for the second group (-4xy+6y^2)/2y -2y(2x-3y) Put it together 3x(2x-3y) -2y(2x-3y) (3x-2y)(2x-3y)